### v1

# NAVINACT: Combining Navigation and Imitation Learning for Bootstrapping Reinforcement Learning

Abstract
Reinforcement Learning (RL) has shown remarkable progress in simulation environments, yet its application to real-world robotic tasks remains limited due to challenges in exploration and generalization. To address these issues, we introduce NAVINACT, a framework that chooses when the robot should use classical motion planning based navigation and when it should learn a policy. To further improve the efficiency in exploration, we use imitation data to bootstrap the exploration. NAVINACT dynamically switches between two modes of operation: navigating to a waypoint using classical techniques when away from the objects and reinforcement learning for fine-grained manipulation control when about to interact with objects. NAVINACT consists of a multi-head architecture composed of ModeNet for mode classification, NavNet for waypoint prediction, and InteractNet for precise manipulation. By combining the strengths of RL and Imitation Learning (IL), NAVINACT improves sample efficiency and mitigates distribution shift, ensuring robust task execution. We evaluate our approach across multiple challenging simulation environments and real-world tasks, demonstrating superior performance in terms of adaptability, efficiency, and generalization compared to existing methods.In both simulated and real-world settings, NAVINACT demonstrates robust performance. In simulations, NAVINACT surpasses baseline methods by 10-15% in training success rates at 30k samples and by 30-40% during evaluation phases. In real-world scenarios, it demonstrates a 30-40% higher success rate on simpler tasks compared to baselines and uniquely succeeds in complex, two-stage manipulation tasks. Datasets and supplementary materials can be found on our website.
keywords:
Reinforcement Learning, Imitation Learning, Efficiency, Generalization, Manipulation
1 Introduction
Figure 1: Multi-headed architecture of NAVINACT: During training, NAVINACT learns to predict waypoints, low-level actions, and the operational mode at each time step. One network (InteractNet) predicts the low-level action $a_{t}$ and the other network (ModeNet) predicts mode $m_{t}$. A separate network (NavNet) predicts the high-level waypoint $w_{t}$. At test time, the system samples $m_{t}$ and either navigates to a waypoint (when $m_{t}=0$) using the predicted waypoint or follows a dense action (when $m_{t}=1$). The architecture allows for dynamic switching between navigation and interaction modes, facilitating seamless transitions and robust performance in complex tasks. An example of how navigation and interaction modes are integrated during execution is shown on the right.
In recent years, reinforcement learning (RL) has made significant strides, achieving remarkable success across various domains [21, 22, 23]. Despite advances in RL, learning behaviors efficiently that generalize to new settings is still a challenge. RL methods are often constrained to short-horizon tasks due to the inherent challenges of exploration in high-dimensional, continuous action spaces, which are characteristic of robotic applications. In these scenarios, reward signals tend to be sparse, making it difficult for a randomly initialized policy to receive any learning signals, especially when the task’s completion is far from the initial state. Additionally, RL typically requires a large number of samples to converge, a requirement that is impractical in real-world environments where large-scale parallel simulation is not feasible. The exploration challenges are further exacerbated when tasks demand both high-level strategic planning and precise low-level control, making it difficult for a single RL model to effectively learn and execute the nuanced decompositions needed for complex manipulation tasks.
To enhance sample efficiency and generalizability, researchers have explored imitation learning (IL) as a means to accelerate learning by leveraging demonstration data. In hierarchical frameworks like LAVA [5], IL is used to parameterize low-level policies based on high-level visual inputs, enabling the system to effectively handle varying types and positions of food during robotic assisted feeding tasks. However, IL alone can lead to distributional shift issues when applied outside the training context. Our focus in this paper is to improve the sample complexity and generalizability of robot learning specifically for manipulation tasks. Our key idea is to improve the exploration process relying on classical motion planning algorithms and imitation data. Specifically, we use motion planning when the robot is “away” from objects and applying RL only when the robot is “near” objects during interactions that require fine control. This strategy reduces the complexity of the learning problem and enhances sample efficiency. Further improvements in sample efficiency can be achieved through imitation learning (IL), though IL alone can raise generalizability issues, which we address by bootstrapping IL with RL.
Our novel approach leverages the hierarchical nature of robotic tasks by implementing a dynamic policy framework that switches between navigating to high-level waypoints and engaging in fine-grained manipulation. We build on HYDRA [29] which introduced the idea of switching between sparse (classical motion planning) and dense (learning) modes. However, the dense mode in HYDRA uses IL which is prone to distributional shift near objects, our method integrates IL with RL, mitigating these shifts and improving task robustness.
While RL is effective in optimizing policies through exploration, it struggles with sparse and delayed rewards. Conversely, IL can expedite learning by providing optimal or near-optimal actions, but it is constrained by static training datasets and the risk of distributional shift when applied to novel contexts. To address these limitations, we combine the strengths of both RL and IL. Methods such as Imitation Bootstrapped Reinforcement Learning (IBRL) [28], [6] and Reinforcement Learning from Prior Data (RLPD) [31] demonstrate this integration, achieving better results compared to using IL or RL alone. However, these methods still require many interaction steps to converge. By incorporating IBRL with mode prediction through ModeNet and waypoint prediction through NavNet, our approach achieves results much faster, significantly improving sample efficiency and performance in complex robotic tasks.
We introduce NAVINACT (NAVigation and INterACTion), a framework that integrates navigation and imitation learning to bootstrap reinforcement learning. Our approach comprises three key components: NavNet for strategic waypoint planning based on visual inputs; ModeNet that dynamically switches between navigation and interaction modes; and InteractNet that performs fine-grained manipulation tasks with precision as shown in Fig. 5. The key contributions of this paper are:
  • We present NAVINACT, a hierarchical policy framework designed to integrate navigation and interaction for complex robotic manipulation tasks.
  • Our method demonstrates adaptability and robustness across various challenging environments, effectively handling both high-level planning and detailed execution.
  • We introduce a generalized mode prediction and waypoint prediction network, reducing the need for task-specific mode labeling and enhancing the system’s scalability.
  • We thoroughly tested NAVINACT in both simulated and real-world settings using a UR3e robot equipped with a Robotiq Hand-e gripper and Realsense cameras. The framework achieved impressive success rates: 85±5% during simulation training and 80±5% in evaluations with unseen environments. In real-world tests, NAVINACT achieved up to 90% success on simpler tasks and 75% on complex two-stage tasks within 30k samples, highlighting its efficiency and adaptability.
2 Related Work
This section explores substantial contributions in reinforcement learning (RL) aimed at enhancing sample efficiency and generalizability, with a focus on methods integrating human demonstrations and strategies to counter distribution shifts in imitation learning (IL). We also examine how leveraging reference policies can optimize RL outcomes.
2.1 Incorporating Model Priors and Refining Action Spaces
Research has increasingly concentrated on embedding structural priors within models to stabilize performance across variable conditions. Object-centric representations and pretrained state models from diverse task datasets have proven to enhance generalization and efficiency [1, 2, 3]. Approaches like [29] HYDRA have refined action spaces to align with these insights. However, NAVINACT goes a step further by dynamically switching action modes, integrating both high-level strategic planning and precise low-level task executions. This not only enhances navigation and manipulation capabilities in intricate settings but also addresses the challenges of generalization and extensive data requirements found in temporal action abstractions and parameterized motion primitives [4, 7, 8, 9, 11, 12]. Building on the foundations laid by previous models, NAVINACT effectively minimizes errors and advances the capabilities of action representation in complex environments, making it a significant advancement in both model integration and action space refinement.
2.2 Planning-Driven Reinforcement Learning
Integrating motion planning with RL has been explored to combine the benefits of both paradigms [13, 14, 15, 16, 17, 18, 19]. For example, GUAPO [13] focuses on single-stage tasks, maintaining RL agents within low pose-estimator uncertainty areas, while Plan-Seq-Learn decomposes complex tasks into sequences of subtasks guided by large language models (LLMs). However, these methods often come with high inference costs or inefficiencies in task decomposition. Our NAVINACT method contrasts these approaches by offering an end-to-end learning solution that is not only faster but also more adaptable to real-time applications, with the potential for further enhancement through LLM integration.
2.3 Reinforcement Learning with Prior Demonstrations
Sample-efficient RL often struggles in sparse reward environments, where initial learning cues are vital. Integrating prior data or human demonstrations into the replay buffer—an approach employed in methods like Reinforcement Learning from Prior Data (RLPD)[31] and IBRL [28]—has proven effective. These methods enhance RL algorithms by oversampling demonstrations during training, significantly improving performance in continuous control domains with techniques like normalization and Q-ensembling [2, 28]. Our NAVINACT framework builds on these principles by dynamically adjusting its strategy based on environmental cues and task demands, using a multi-headed architecture that optimizes both navigation and interaction. This approach not only leverages prior demonstrations for improved initial guidance but also adapts in real-time to enhance task execution and generalization in complex robotic manipulation tasks.
2.4 Advancing RL Sample Efficiency
Recent advancements in RL have prioritized enhancing sample efficiency through innovative regularization strategies. Techniques like RED-Q [20] and DropoutQ [24] apply regularization to the Q-function, utilizing ensembling or dropout to achieve higher update-to-data (UTD) ratios and expedited convergence. These approaches prove particularly effective in state-based RL, while in pixel-based settings, strategies such as image augmentation with random shifts [25] boost performance without additional computational demands. Building upon these concepts, the IBRL [28] framework integrates imitation learning (IL) with RL, harnessing a pre-trained IL policy to direct the RL training, enhancing the initial learning curve. However, IBRL still necessitates frequent updates to achieve optimal performance levels. In contrast, our NAVINACT method refines this approach by significantly reducing the number of updates required, effectively merging the benefits of IL, RL as well as mode switching and navigation to heighten sample efficiency and streamline learning processes.
3 NAVINACT
We consider a standard Markov decision process (MDP) consisting of state space $s\in\mathcal{S}$, continuous action space $\mathcal{A}=[-1,1]^{d}$, deterministic state transition function $T:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$, sparse reward function $R:\mathcal{S}\times\mathcal{A}\to\{0,1\}$ that returns 1 when the task is completed and 0 otherwise, and discount factor $\gamma$.
Figure 2: ModeNet architecture for classifying modes (navigate vs. interact) based on visual inputs.
Figure 3: NavNet architecture for predicting waypoints to guide high-level navigation tasks.
Our proposed method, NAVINACT, features a multi-headed architecture designed for efficient robotic manipulation. The system is composed of three main components: ModeNet, NavNet, and InteractNet, each contributing to different levels of task execution. The overall architecture, as illustrated in Fig. 1, includes heads $\pi^{M}_{\theta}:\mathcal{S}\rightarrow\{0,1\}$, $\pi^{W}_{\theta}:\mathcal{S}\rightarrow P$, and $\pi^{A}_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$, corresponding to mode classification, high-level waypoint prediction, and low-level action prediction respectively, where $P$ is robot proprioceptive data such as end effector poses and gripper widths. This design enables dynamic switching between coarse and fine-grained control during task execution.
3.1 ModeNet: A Vision-Based Mode Classification Network
ModeNet is a deep learning architecture that classifies the operational mode of the robot based on visual input. It determines whether the robot should operate in navigation mode, or in interaction mode, as shown in Fig. 5.
ModeNet utilizes a ResNet backbone as shown in Fig. 3. The architecture consists of a feature extractor followed by a mode classifier. The feature extractor employs a pretrained ResNet model to capture high-level spatial and contextual features from input images, while the mode classifier processes these features to predict the robot’s operational mode.
3.2 NavNet: A Vision-Based Waypoint Prediction Network
NavNet is designed to predict waypoints for robotic manipulation tasks by processing both visual and proprioceptive inputs. It generates high-level action outputs that guide the robot towards the target objects in the environment. NavNet combines a convolutional neural network (CNN) for visual processing with fully connected layers that integrate visual features with proprioceptive data, Fig. 3. This design allows for efficient waypoint prediction, facilitating robust navigation in complex environments. Once we have the predicted waypoint, we use a sampling-based motion planner, AIT* [39], to perform motion planning.
Algorithm 1 NAVINACT. Major modifications with framework highlighted in blue.
1:  Hyperparameters: Number of critics $E$, number of critic updates $G$, update frequency $U$, exploration std $\sigma$, noise clip $c$, Complexity Function $g$
2:  Train imitation policy $\mu_{\psi}$ on demonstrations $D=\{\xi_{1},\dots,\xi_{n}\}$ using the selected IL algorithm.
3:  Train ModeNet on dataset $D_{\text{mode}}$ to classify mode (navigate vs interact).
4:  Train NavNet on dataset $D_{\text{nav}}$ to predict waypoints for navigation tasks.
5:  Initialize policy $\pi_{\theta}$, target policy $\pi_{\theta^{\prime}}$, and critics $Q_{\phi}$, target critics $Q_{\phi^{\prime}}$ for $i=1,\dots,E$
6:  Initialize replay buffer $B$ with demonstrations $\{\xi_{1},\dots,\xi_{n}\}$
7:  for $t=1$ to $N$ do
8:     Observe current state $s_{t}$ from the environment
9:     Determine mode $m_{t}$ using ModeNet.
10:     if mode $m_{t}$ is “navigate” then
11:        Predict waypoint $w_{t}$ using NavNet
12:        Generate navigation action $a^{\text{Nav}}_{t}$ based on waypoint $w_{t}$
13:        Execute navigation action $a^{\text{Nav}}_{t}$
14:     else
15:        Compute IL action $a^{\text{IL}}_{t}\sim\mu_{\psi}(s_{t})$ and RL action $a^{\text{RL}}_{t}=\pi_{\theta}(s_{t})+\epsilon$, where $\epsilon\sim N(0,\sigma^{2})$
16:        Sample a set $K$ of 2 indices from $\{1,2,\dots,E\}$
17:        Select action $a_{t}$ with higher Q-value from $\{a^{\text{IL}},a^{\text{RL}}\}$
18:        Execute action $a_{t}$
19:        Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in replay buffer $B$
20:        if $t\%U=0$ then
21:           Perform TD3 update using minibatches from replay buffer $B$ [38]
22:        end if
23:     end if
24:  end for
3.3 InteractNet: Low-Level Action Prediction
InteractNet integrates imitation learning (IL) with reinforcement learning (RL) to enhance sample efficiency in robotic task execution. Initially, it trains using expert demonstrations to establish a baseline IL policy. This IL policy guides the RL training phase by recommending high-quality actions, aiding in effective learning through techniques like Temporal Difference (TD) and off-policy RL methods such as TD3. Actions are selected based on the highest Q-values, comparing inputs from both IL and RL strategies.
Figure 4: Analysis of action selection from Behavior Cloning (BC) policy during NAVINACT training. The figures show the proportion of actions taken from the BC and RL policy across different tasks, providing insight into the system’s reliance on BC during training.
Fig. 4 illustrates how often IL actions are chosen throughout training. Early on, the predominance of RL actions is noticeable due to initial high “fake” Q-values. As the system updates and learns from real performance data, IL actions become more frequent, underscoring their effectiveness. Over time, while the reliance on IL decreases as the RL policy matures, IL actions remain crucial, particularly in complex tasks where RL alone is insufficient for full policy convergence. This ongoing use of IL actions at all stages indicates their integral role in ensuring robust and efficient decision-making within NAVINACT.
4 Experiments
In this section, we present a comprehensive evaluation of NAVINACT, focusing on its sample efficiency, generalizability, and performance in real-world scenarios. We conducted experiments in three challenging simulation environments (Fig. 8) and two complex real-world tasks (Fig. 10).
4.1 Training and evaluation of ModeNet and NavNet
Figure 5: Sequence of images illustrating ModeNet’s predicted modes during trajectory execution.
(a) Assembly Env
(b) BoxClose Env
(c) Coffeepush Env
Figure 6: This figure displays the evaluation setups for three MetaWorld tasks: Assembly, BoxClose, and CoffeePush, with randomized initial gripper positions. Each subfigure illustrates how the gripper’s starting position varies in these environments, testing NAVINACT’s ability to adapt to different initial conditions.
Can we predict the modes and waypoints accurately?
ModeNet is trained on a dataset comprising 1,500 environmental camera images for simulation. The ground truth labels were determined by a distance threshold between the end-effector and the object, effectively deciding whether the robot should be in a navigation mode (moving towards an object) or an interaction mode (engaging directly with an object). Data augmentation techniques like rotation and contrast adjustment were applied during training. The network utilizes cross-entropy loss and is optimized with Adam. The network’s performance is evaluated using metrics such as accuracy, precision, recall, and F1 score. ModeNet achieved an accuracy score of 0.89, an F1 score of 0.85, a recall score of 0.9, and a precision score of 0.8, demonstrating its effectiveness in accurately classifying operational modes for robotic tasks.
Figure 7: Sequence showing ModeNet predicting mode 0, followed by NavNet guiding the robotic arm to the designated waypoint.
4.1.1 NavNet
NavNet is trained on the same dataset of 1,500 environmental camera images as ModeNet. The ground truth for NavNet involves waypoints defined as positions 2 cm above the objects, guiding the robot on where to navigate next. Like ModeNet, data augmentation techniques such as rotation and contrast adjustments were used to enhance training robustness. NavNet employs Mean Squared Error (MSE) loss and is optimized with the Adam optimizer. It achieved an impressive accuracy score of 0.93, indicating high effectiveness in predicting accurate navigation waypoints based on visual inputs. Fig. 7 shows basic demo of NAVINACT in navigation mode.
4.1.2 ModeNet and NavNet for realworld experiments
For real-world experiments, ModeNet and NavNet were trained using a more extensive dataset, incorporating both wrist and environment camera images totaling 6,200. This approach enhanced the diversity and representativeness of the training data, ensuring robustness and accuracy in real-world applications.
4.2 Simulation Experiments
Our evaluation suite includes three tasks from MetaWorld [37], each using sparse 0/1 task completion rewards at the end of each episode. These tasks are a subset of those evaluated in IBRL [28] and cover medium and hard difficulty levels as categorized in [35].
The specific environments used are assembly-v2, box-close-v2, and coffee-push-v2. Given the absence of human demonstrations in MetaWorld, we utilized the demonstration datasets generated by IBRL, which were created using scripted expert policies from [37]. Although these tasks are relatively simple, they are sufficient to differentiate between stronger methods, especially in terms of sample efficiency and generalizability.
Implementation of NAVINACT and Baselines:
NAVINACT employs Behavior Cloning (BC) for imitation learning (IL) and Twin Delayed Deep Deterministic Policy Gradient (TD3) for reinforcement learning (RL). A ResNet-18 vision encoder is used for BC, and random-shift image augmentation is applied to enhance RL performance. NAVINACT is benchmarked against three baselines: IBRL [28], RL-MN (RL with ModeNet and NavNet), and standard RL.
IBRL uses a pre-trained IL policy, $\mu_{\psi}$, to aid exploration during RL and to assist in target value estimation during Temporal Difference (TD) learning. Likewise, NAVINACT also uses IL policy within the InteractNet, that speeds up RL during object interaction. RL with ModeNet and NavNet, which we refer as RL-MN, incorporates ModeNet and NavNet to accelerate exploration by predicting waypoints, reducing the number of exploration steps needed in RL. Standard RL utilizes TD3 but does not use any IL or hierarchical mode prediction components.
4.3 Results
Following the training of NAVINACT and the baselines, we evaluate its performance on unseen environments and compare it against baseline methods. Performance is assessed using training and evaluation scores, which indicate sample efficiency and generalizability. Fig. 8 illustrates the comparative results for the all the metaworld environment, including training and evaluation scores for various initial positions.
(a) Assembly Env
(b) Boxclose Env
(c) CoffeePush Env
(d) Training with random inital Objects position.
(e) Training with random inital Objects position.
(f) Training with random inital Objects position.
(g) Evaluation with random initial Objects position.
(h) Evaluation with random initial Objects position.
(i) Evaluation with random initial Objects position.
(j) Evaluation with random initial Objects and gripper position.
(k) Evaluation with random initial Objects and gripper position.
(l) Evaluation with random initial Objects and gripper position.
Figure 8: Performance Evaluation of NAVINACT Across MetaWorld Tasks. This figure presents the training and evaluation results for three challenging MetaWorld environments: Assembly, BoxClose, and CoffeePush. The top row depicts the environments in their initial configurations. The middle row shows the training performance of NAVINACT under conditions where object positions are randomized. The bottom two rows illustrate the evaluation performance under different randomized scenarios: varying initial object positions and combined randomization of both object and gripper positions. The results demonstrate NAVINACT’s robustness and adaptability to varying task complexities, maintaining high performance across different training and evaluation conditions.
Both NAVINACT and IBRL utilize a trained BC policy, with NAVINACT achieving a higher evaluation score of 0.38 in the assembly environment, while RL-MN and RL do not incorporate this policy. Our evaluation is focused on 3 key metrics — training score for the simulation tasks with randomization with respect to objects position, evaluation score with randomization with respect to objects position and evaluation score with randomization with respect to random initial gripper (as shown in Fig. 6 as well as object positions. The performance of NAVINACT, IBRL, RL-MN, and standard RL is compared across the Meta-World environments over 60k timesteps and 7 seeds.
NAVINACT outperforms baselines across all tasks in terms of both sample efficiency and final performance, solving all tasks within 30K samples.
4.3.1 Does NAVINACT Lead to Better Sample Efficiency?
To assess sample efficiency, we compare the training curves of NAVINACT with those of IBRL, RL-MN, and RL across various MetaWorld tasks. These tasks represent different levels of complexity, categorized into medium, hard, and very hard tiers. NAVINACT demonstrates better sample efficiency, as evidenced by faster convergence and higher success rates with fewer interaction steps as shown in Fig. 8(d),8(e) and 8(f). This improvement is attributed to NAVINACT’s mode prediction, which effectively reduces task complexity by narrowing down the action space at each step. NAVINACT consistently achieves an impressive success rate of 85±5% during training and learns quickly within 30k sample steps.
Figure 9: Experimental Setup
4.3.2 Does NAVINACT Generalize to Unseen Conditions?
To evaluate generalization, we test NAVINACT in conditions not included in the training data, such as variations in object positions, and initial gripper position. NAVINACT shows strong generalization capabilities, significantly outperforming IBRL and pure RL methods in unseen scenarios as indicated in Fig. 8(g),8(h),8(i),8(j),8(k) and 8(l). The mode prediction network plays a crucial role in adapting to new conditions by dynamically adjusting the operational mode, which is not achievable by methods solely relying on IL. Our algorithm maintains a nearly equivalent rate of 80±5% in evaluation phases across unseen environments. NAVINACT consistently outperforms other baselines in MetaWorld tasks, demonstrating its robustness to variability and changes in the environment.
We see that, NAVINACT matches or surpasses other baselines in Meta-World. The results indicate that, NAVINACT outperforms the other three methods in both training and evaluation. Notably, as randomization increases, the performance of IBRL, RL-MN, and standard RL decreases, while NAVINACT maintains consistent performance and demonstrates the robustness to variability and changes in environment.
4.4 Real-world Experimental Setup
We design two tasks named Lift and pick-and-place as shown in Fig.10(a), 10(b), 10(c). The tasks use a UR3e robot. The robot is equipped with a Robotiq hand-e gripper and a realsense camera mounted on the wrist. We also use another realsense camera as environment camera for the experiment as shown in Fig. 9.
Actions are 4-dimensional consisting of 3 dimensions for difference of end-effector position (deltas), under a Cartesian Delta controller and 1 dimension for absolute position of the gripper. Policies run at 6 Hz. For each task, we collect a small number of prior demonstrations via teleoperation with a passive arm Gello [36], that we modified for UR3e controller, and then run different RL methods for a fixed number of interaction steps. All methods use the exact same hyper-parameters and network architectures as in metaworld tasks. We illustrate these tasks in Fig. 10 and briefly describe them here.
Lift: The objective is to pick up a foam block. The initial location of the block is randomized over roughly $22cm\times(20cm-25cm)$ trapezoid, which covers the entire area visible from the wrist-camera when the robot is at the home position. See Fig.10(a), Fig. 10(b) for reference. We collect 10 demonstrations for this task due to its simplicity. It uses wrist-camera images as observations. We detect whether the gripper is holding the block by checking if the gripper width is static and the desired gripper width is smaller than the actual gripper width. The success detector returns 1 if the end effector has move upward by at least 2cm while holding the block. The maximum episode length is 120.
Pick-and-Place: The objective is to pick a soft toy on a wooden block. The initial location of the soft toy and the block is in a fixed position. The soft toy is initialized so that its visible from the wrsit camera view, see Fig.10(c). We use 20 prior demonstrations. This task uses environment camera images as observations as well because the wrist-camera loses sight of the block after picking up the soft-toy. We detect whether the gripper is holding the oft-toy by checking if the gripper width is static and the desired gripper width is smaller than the actual gripper width. The success returns 1 when the gripper width is static, the end-effector position are near the threshold of desired block position. The maximum episode legnth is 300.
(a) Lift
(b) Lift with randomization
(c) Pick and Place
Figure 10: Top: Illustrations of lift task and the variation in the initialization of the task, left (fixed) right (randomized block position). Bottom: Performance evaluation on the lift task under both fixed and randomized conditions. The training curves depict the success rate over time, demonstrating NAVINACT’s effective learning and adaptability in varying scenarios.
As the primary goal of our real-world evaluations is to compare sample-efficiency and performance of various algorithms, we design rule-based success detectors and perform manual reset between episodes to ensure accurate reward and initial conditions. Note that sparse 0/1 reward from the success detector is the only source of reward.
4.5 Real-world Experiment Results
Does NAVINACT Work Effectively in Real-World Scenarios? Fig. 10 presents the training curves of NAVINACT compared to IBRL and across all different tasks. Each task is allowed a different interaction budget based on its difficulty, and the success rate is measured over each 1000-step interval during training. NAVINACT consistently learns faster than IBRL, and in the most challenging task, Pick-and-Place, it is the only method that surpasses behavior cloning (BC) under a 5K interaction budget.
In real-world evaluations, we assessed the final performance of each method using identical initial conditions. In the Lift task, NAVINACT achieves the highest success rate in both regular and “Hard Eval” settings, where the block’s initial position is more challenging. Even when the task difficulty increases, NAVINACT maintains a high success of 90% rate, demonstrating its robustness to distributional shifts and unseen scenarios, unlike BC, which fails in harder settings due to its reliance on the static training dataset as shown in Fig. 10.
How does NAVINACT perform for long-horizon tasks? The Pick-and-Place task, the most difficult due to two stage task manipulation, further highlights NAVINACT’s advantages as displayed in Fig. 10. While BC also doesn’t performs well due to suboptimal demostration, other method IBRL fail to match its success rate. While BC achieves a limited success rate of 20%, NAVINACT surpasses this by reaching a 75% success rate, even in complex scenarios where small errors can lead to significant failures.
5 Discussion
5.1 Conclusion
In this paper, we presented NAVINACT, a hierarchical policy framework that combines NavNet for waypoint planning, ModeNet for dynamic mode switching, and InteractNet for fine-grained manipulation. By integrating imitation learning (IL) with reinforcement learning (RL), NAVINACT effectively addresses challenges of sample efficiency and distributional shift, outperforming existing methods. Through extensive simulations and real-world experiments, we demonstrated NAVINACT’s superior performance, generalizability, and adaptability across various tasks. Our approach learns faster and requires fewer interaction steps due to the innovative integration of mode and waypoint prediction, making it a scalable and robust solution for complex robotic manipulation in real-world scenarios.
5.2 Limitation and Future Work
While NAVINACT demonstrates significant advancements in sample efficiency and generalizability, one limitation of our approach is the reliance on collecting high-quality demonstration data, which can be time-consuming and expensive. This dependence on expert-generated data constrains the scalability and applicability of our framework across a wider range of tasks and environments. Future work could focus on developing methods that allow non-experts to collect usable data quickly and efficiently, potentially through resources such as videos in the wild.

### v2

@@
-NAVINACT: Combining Navigation and Imitation Learning for Bootstrapping Reinforcement Learning
+PLANRL: A Motion Planning and Imitation Learning Framework to Bootstrap Reinforcement Learning
 Abstract
 Reinforcement Learning (RL) has shown remarkable progress in simulation environments, yet its application to real-world robotic tasks remains limited due to challenges in exploration and generalization.
-To address these issues, we introduce NAVINACT, a framework that chooses when the robot should use classical motion planning based navigation and when it should learn a policy.
+To address these issues, we introduce PLANRL, a framework that chooses when the robot should use classical motion planning and when it should learn a policy.
 To further improve the efficiency in exploration, we use imitation data to bootstrap the exploration.
-NAVINACT dynamically switches between two modes of operation: navigating to a waypoint using classical techniques when away from the objects and reinforcement learning for fine-grained manipulation control when about to interact with objects.
-NAVINACT consists of a multi-head architecture composed of ModeNet for mode classification, NavNet for waypoint prediction, and InteractNet for precise manipulation.
-By combining the strengths of RL and Imitation Learning (IL), NAVINACT improves sample efficiency and mitigates distribution shift, ensuring robust task execution.
-We evaluate our approach across multiple challenging simulation environments and real-world tasks, demonstrating superior performance in terms of adaptability, efficiency, and generalization compared to existing methods.In both simulated and real-world settings, NAVINACT demonstrates robust performance.
-In simulations, NAVINACT surpasses baseline methods by 10-15% in training success rates at 30k samples and by 30-40% during evaluation phases.
+PLANRL dynamically switches between two modes of operation: reaching a waypoint using classical techniques when away from the objects and reinforcement learning for fine-grained manipulation control when about to interact with objects.
+PLANRL architecture is composed of ModeNet for mode classification, NavNet for waypoint prediction, and InteractNet for precise manipulation.
+By combining the strengths of RL and Imitation Learning (IL), PLANRL improves sample efficiency and mitigates distribution shift, ensuring robust task execution.
+We evaluate our approach across multiple challenging simulation environments and real-world tasks, demonstrating superior performance in terms of adaptability, efficiency, and generalization compared to existing methods.
+In simulations, PLANRL surpasses baseline methods by 10-15% in training success rates at 30k samples and by 30-40% during evaluation phases.
 In real-world scenarios, it demonstrates a 30-40% higher success rate on simpler tasks compared to baselines and uniquely succeeds in complex, two-stage manipulation tasks.
@@
 1 Introduction
-Figure 1: Multi-headed architecture of NAVINACT: During training, NAVINACT learns to predict waypoints, low-level actions, and the operational mode at each time step.
+Figure 1: Architecture of PLANRL: During training, PLANRL learns to predict waypoints, low-level actions, and the operational mode at each time step.
 One network (InteractNet) predicts the low-level action $a_{t}$ and the other network (ModeNet) predicts mode $m_{t}$.
 A separate network (NavNet) predicts the high-level waypoint $w_{t}$.
-At test time, the system samples $m_{t}$ and either navigates to a waypoint (when $m_{t}=0$) using the predicted waypoint or follows a dense action (when $m_{t}=1$).
-The architecture allows for dynamic switching between navigation and interaction modes, facilitating seamless transitions and robust performance in complex tasks.
-An example of how navigation and interaction modes are integrated during execution is shown on the right.
+At test time, the system samples $m_{t}$ and either moves to a waypoint (when $m_{t}=0$) using the predicted waypoint or follows a dense action (when $m_{t}=1$).
+The architecture allows for dynamic switching between motion-planning and interaction modes, facilitating robust performance in complex tasks.
+An example of how motion planning and interaction modes are integrated during execution is shown on the right.
 In recent years, reinforcement learning (RL) has made significant strides, achieving remarkable success across various domains [21, 22, 23].
@@
 To enhance sample efficiency and generalizability, researchers have explored imitation learning (IL) as a means to accelerate learning by leveraging demonstration data.
-In hierarchical frameworks like LAVA [5], IL is used to parameterize low-level policies based on high-level visual inputs, enabling the system to effectively handle varying types and positions of food during robotic assisted feeding tasks.
+In hierarchical frameworks like LAVA [5], IL is used to parameterize low-level policies based on high-level visual inputs, enabling the system to effectively handle varying types and positions of food during robotic-assisted feeding tasks.
 However, IL alone can lead to distributional shift issues when applied outside the training context.
@@
 Our key idea is to improve the exploration process relying on classical motion planning algorithms and imitation data.
-Specifically, we use motion planning when the robot is “away” from objects and applying RL only when the robot is “near” objects during interactions that require fine control.
+Specifically, we use motion planning when the robot is “away” from objects and apply RL only when the robot is “near” objects during interactions that require fine control.
 This strategy reduces the complexity of the learning problem and enhances sample efficiency.
 Further improvements in sample efficiency can be achieved through imitation learning (IL), though IL alone can raise generalizability issues, which we address by bootstrapping IL with RL.
-Our novel approach leverages the hierarchical nature of robotic tasks by implementing a dynamic policy framework that switches between navigating to high-level waypoints and engaging in fine-grained manipulation.
+Our novel approach leverages the hierarchical nature of robotic tasks by implementing a dynamic policy framework that switches between moving to high-level waypoints and engaging in fine-grained manipulation.
 We build on HYDRA [29] which introduced the idea of switching between sparse (classical motion planning) and dense (learning) modes.
@@
 By incorporating IBRL with mode prediction through ModeNet and waypoint prediction through NavNet, our approach achieves results much faster, significantly improving sample efficiency and performance in complex robotic tasks.
-We introduce NAVINACT (NAVigation and INterACTion), a framework that integrates navigation and imitation learning to bootstrap reinforcement learning.
-Our approach comprises three key components: NavNet for strategic waypoint planning based on visual inputs; ModeNet that dynamically switches between navigation and interaction modes; and InteractNet that performs fine-grained manipulation tasks with precision as shown in Fig.
+We introduce PLANRL (NAVigation and INterACTion), a framework that integrates motion-planning and imitation learning to bootstrap reinforcement learning.
+Our approach comprises three key components: NavNet for strategic waypoint planning based on visual inputs; ModeNet that dynamically switches between motion-planning and interaction modes; and InteractNet that performs fine-grained manipulation tasks with precision as shown in Fig.
 5.
 The key contributions of this paper are:
-  • We present NAVINACT, a hierarchical policy framework designed to integrate navigation and interaction for complex robotic manipulation tasks.
+  • We present PLANRL, a hierarchical policy framework designed to integrate motion-planning and interaction for complex robotic manipulation tasks.
   • Our method demonstrates adaptability and robustness across various challenging environments, effectively handling both high-level planning and detailed execution.
   • We introduce a generalized mode prediction and waypoint prediction network, reducing the need for task-specific mode labeling and enhancing the system’s scalability.
-  • We thoroughly tested NAVINACT in both simulated and real-world settings using a UR3e robot equipped with a Robotiq Hand-e gripper and Realsense cameras.
+  • We thoroughly tested PLANRL in both simulated and real-world settings using a UR3e robot equipped with a Robotiq Hand-e gripper and Realsense cameras.
 The framework achieved impressive success rates: 85±5% during simulation training and 80±5% in evaluations with unseen environments.
-In real-world tests, NAVINACT achieved up to 90% success on simpler tasks and 75% on complex two-stage tasks within 30k samples, highlighting its efficiency and adaptability.
+In real-world tests, PLANRL achieved up to 90% success on simpler tasks and 75% on complex two-stage tasks within 30k samples, highlighting its efficiency and adaptability.
 2 Related Work
@@
 Approaches like [29] HYDRA have refined action spaces to align with these insights.
-However, NAVINACT goes a step further by dynamically switching action modes, integrating both high-level strategic planning and precise low-level task executions.
-This not only enhances navigation and manipulation capabilities in intricate settings but also addresses the challenges of generalization and extensive data requirements found in temporal action abstractions and parameterized motion primitives [4, 7, 8, 9, 11, 12].
-Building on the foundations laid by previous models, NAVINACT effectively minimizes errors and advances the capabilities of action representation in complex environments, making it a significant advancement in both model integration and action space refinement.
+However, PLANRL goes a step further by dynamically switching action modes, integrating both high-level strategic planning and precise low-level task executions.
+This not only enhances motion-planning and manipulation capabilities in intricate settings but also addresses the challenges of generalization and extensive data requirements found in temporal action abstractions and parameterized motion primitives [4, 7, 8, 9, 11, 12].
+Building on the foundations laid by previous models, PLANRL effectively minimizes errors and advances the capabilities of action representation in complex environments, making it a significant advancement in both model integration and action space refinement.
 2.2 Planning-Driven Reinforcement Learning
 Integrating motion planning with RL has been explored to combine the benefits of both paradigms [13, 14, 15, 16, 17, 18, 19].
-For example, GUAPO [13] focuses on single-stage tasks, maintaining RL agents within low pose-estimator uncertainty areas, while Plan-Seq-Learn decomposes complex tasks into sequences of subtasks guided by large language models (LLMs).
+For example, GUAPO [13] focuses on single-stage tasks, maintaining RL agents within low pose-estimator uncertainty areas, while Plan-Seq-Learn [30] decomposes complex tasks into sequences of subtasks guided by large language models (LLMs).
 However, these methods often come with high inference costs or inefficiencies in task decomposition.
-Our NAVINACT method contrasts these approaches by offering an end-to-end learning solution that is not only faster but also more adaptable to real-time applications, with the potential for further enhancement through LLM integration.
+Our PLANRL method contrasts these approaches by offering an end-to-end learning solution that is not only faster but also more adaptable to real-time applications, with the potential for further enhancement through LLM integration.
 2.3 Reinforcement Learning with Prior Demonstrations
@@
 These methods enhance RL algorithms by oversampling demonstrations during training, significantly improving performance in continuous control domains with techniques like normalization and Q-ensembling [2, 28].
-Our NAVINACT framework builds on these principles by dynamically adjusting its strategy based on environmental cues and task demands, using a multi-headed architecture that optimizes both navigation and interaction.
+Our PLANRL framework builds on these principles by dynamically adjusting its strategy based on environmental cues and task demands, using an architecture that optimizes both motion-planning and interaction.
 This approach not only leverages prior demonstrations for improved initial guidance but also adapts in real-time to enhance task execution and generalization in complex robotic manipulation tasks.
@@
 These approaches prove particularly effective in state-based RL, while in pixel-based settings, strategies such as image augmentation with random shifts [25] boost performance without additional computational demands.
-Building upon these concepts, the IBRL [28] framework integrates imitation learning (IL) with RL, harnessing a pre-trained IL policy to direct the RL training, enhancing the initial learning curve.
+Building upon these concepts, the IBRL [28] framework combines imitation learning (IL) with reinforcement learning (RL) by leveraging a pre-trained IL policy to guide RL training from the start, allowing the RL agent to benefit from high-quality actions provided by the IL policy.
+This integration significantly accelerates both exploration and training efficiency.
 However, IBRL still necessitates frequent updates to achieve optimal performance levels.
-In contrast, our NAVINACT method refines this approach by significantly reducing the number of updates required, effectively merging the benefits of IL, RL as well as mode switching and navigation to heighten sample efficiency and streamline learning processes.
-3 NAVINACT
+In contrast, our PLANRL method refines this approach by significantly reducing the number of updates required, effectively merging the benefits of IL, RL as well as mode switching and motion-planning to heighten sample efficiency and streamline learning processes.
+3 PLANRL
 We consider a standard Markov decision process (MDP) consisting of state space $s\in\mathcal{S}$, continuous action space $\mathcal{A}=[-1,1]^{d}$, deterministic state transition function $T:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$, sparse reward function $R:\mathcal{S}\times\mathcal{A}\to\{0,1\}$ that returns 1 when the task is completed and 0 otherwise, and discount factor $\gamma$.
-Figure 2: ModeNet architecture for classifying modes (navigate vs. interact) based on visual inputs.
-Figure 3: NavNet architecture for predicting waypoints to guide high-level navigation tasks.
-Our proposed method, NAVINACT, features a multi-headed architecture designed for efficient robotic manipulation.
+Figure 2: ModeNet architecture designed to classify modes (motion-planning vs. interact) based on visual inputs.
+Figure 3: NavNet architecture for predicting waypoints to guide high-level motion-planning tasks.
+Our proposed method, PLANRL, features an architecture designed for efficient robotic manipulation.
 The system is composed of three main components: ModeNet, NavNet, and InteractNet, each contributing to different levels of task execution.
 The overall architecture, as illustrated in Fig.
-1, includes heads $\pi^{M}_{\theta}:\mathcal{S}\rightarrow\{0,1\}$, $\pi^{W}_{\theta}:\mathcal{S}\rightarrow P$, and $\pi^{A}_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$, corresponding to mode classification, high-level waypoint prediction, and low-level action prediction respectively, where $P$ is robot proprioceptive data such as end effector poses and gripper widths.
+1, includes heads $\pi^{M}_{\theta}:\mathcal{S}\rightarrow\{0,1\}$, $\pi^{W}_{\theta}:\mathcal{S}\rightarrow W$, and $\pi^{A}_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$, corresponding to mode classification, high-level waypoint prediction, and low-level action prediction respectively, where $W$ is robot end-effector position.
 This design enables dynamic switching between coarse and fine-grained control during task execution.
@@
 ModeNet is a deep learning architecture that classifies the operational mode of the robot based on visual input.
-It determines whether the robot should operate in navigation mode, or in interaction mode, as shown in Fig.
+It determines whether the robot should operate in motion-planning mode, or in interaction mode, as shown in Fig.
 5.
@@
 The architecture consists of a feature extractor followed by a mode classifier.
-The feature extractor employs a pretrained ResNet model to capture high-level spatial and contextual features from input images, while the mode classifier processes these features to predict the robot’s operational mode.
+The feature extractor employs a pre-trained ResNet model to capture high-level spatial and contextual features from input images, while the mode classifier processes these features to predict the robot’s operational mode.
 3.2 NavNet: A Vision-Based Waypoint Prediction Network
-NavNet is designed to predict waypoints for robotic manipulation tasks by processing both visual and proprioceptive inputs.
-It generates high-level action outputs that guide the robot towards the target objects in the environment.
-NavNet combines a convolutional neural network (CNN) for visual processing with fully connected layers that integrate visual features with proprioceptive data, Fig.
+NavNet is designed to predict waypoints for robotic manipulation tasks by processing visual inputs.
+It generates high-level action outputs that guides the robot towards the target objects in the environment.
+NavNet uses a convolutional neural network (CNN) for visual processing, effectively handling the image data to produce actionable waypoints, Fig.
 3.
-This design allows for efficient waypoint prediction, facilitating robust navigation in complex environments.
-Once we have the predicted waypoint, we use a sampling-based motion planner, AIT* [39], to perform motion planning.
-Algorithm 1 NAVINACT. Major modifications with framework highlighted in blue.
+Once we have the predicted waypoint, we use a sampling-based motion planner, AIT* [39], to perform motion planning as it requires minimal setup—just collision checking—and performs well in planning tasks.
+Algorithm 1 PLANRL. Major modifications with framework highlighted in blue.
 1:  Hyperparameters: Number of critics $E$, number of critic updates $G$, update frequency $U$, exploration std $\sigma$, noise clip $c$, Complexity Function $g$
 2:  Train imitation policy $\mu_{\psi}$ on demonstrations $D=\{\xi_{1},\dots,\xi_{n}\}$ using the selected IL algorithm.
-3:  Train ModeNet on dataset $D_{\text{mode}}$ to classify mode (navigate vs interact).
-4:  Train NavNet on dataset $D_{\text{nav}}$ to predict waypoints for navigation tasks.
+3:  Train ModeNet on dataset $D_{\text{mode}}$ to classify mode (motion-planning vs interaction).
+4:  Train NavNet on dataset $D_{\text{nav}}$ to predict waypoints for motion-planning tasks.
 5:  Initialize policy $\pi_{\theta}$, target policy $\pi_{\theta^{\prime}}$, and critics $Q_{\phi}$, target critics $Q_{\phi^{\prime}}$ for $i=1,\dots,E$
@@
 9:     Determine mode $m_{t}$ using ModeNet.
-10:     if mode $m_{t}$ is “navigate” then
+10:     if mode $m_{t}$ is 0 then
 11:        Predict waypoint $w_{t}$ using NavNet
-12:        Generate navigation action $a^{\text{Nav}}_{t}$ based on waypoint $w_{t}$
-13:        Execute navigation action $a^{\text{Nav}}_{t}$
+12:        Generate motion-planning action $a^{\text{Nav}}_{t}$ based on waypoint $w_{t}$
+13:        Execute motion-planning action $a^{\text{Nav}}_{t}$
 14:     else
@@
 4 illustrates how often IL actions are chosen throughout training.
-Early on, the predominance of RL actions is noticeable due to initial high “fake” Q-values.
-As the system updates and learns from real performance data, IL actions become more frequent, underscoring their effectiveness.
-Over time, while the reliance on IL decreases as the RL policy matures, IL actions remain crucial, particularly in complex tasks where RL alone is insufficient for full policy convergence.
-This ongoing use of IL actions at all stages indicates their integral role in ensuring robust and efficient decision-making within NAVINACT.
+Early on, the predominance of RL actions is noticeable due to initially overestimated Q-values.
+As the system updates and learns from real performance data, IL actions become more frequent, recognizing their effectiveness from expert demonstrations in the replay buffer.
+Over time, although reliance on IL actions decreases as the RL policy matures, they remain essential in complex tasks where RL alone is insufficient for full policy convergence.
+This consistent integration of IL actions at all stages underscores their critical role in ensuring robust and efficient decision-making within PLANRL.
 4 Experiments
-In this section, we present a comprehensive evaluation of NAVINACT, focusing on its sample efficiency, generalizability, and performance in real-world scenarios.
+In this section, we present a comprehensive evaluation of PLANRL, focusing on its sample efficiency, generalizability, and performance in real-world scenarios.
 We conducted experiments in three challenging simulation environments (Fig.
@@
 Figure 6: This figure displays the evaluation setups for three MetaWorld tasks: Assembly, BoxClose, and CoffeePush, with randomized initial gripper positions.
-Each subfigure illustrates how the gripper’s starting position varies in these environments, testing NAVINACT’s ability to adapt to different initial conditions.
+Each subfigure illustrates how the gripper’s starting position varies in these environments, testing PLANRL’s ability to adapt to different initial conditions.
 Can we predict the modes and waypoints accurately?
 ModeNet is trained on a dataset comprising 1,500 environmental camera images for simulation.
-The ground truth labels were determined by a distance threshold between the end-effector and the object, effectively deciding whether the robot should be in a navigation mode (moving towards an object) or an interaction mode (engaging directly with an object).
+The ground truth labels were determined by a distance threshold between the end-effector and the object, effectively deciding whether the robot should be in a motion-planning mode (moving towards an object) or an interaction mode (engaging directly with an object).
 Data augmentation techniques like rotation and contrast adjustment were applied during training.
@@
 4.1.1 NavNet
-NavNet is trained on the same dataset of 1,500 environmental camera images as ModeNet.
-The ground truth for NavNet involves waypoints defined as positions 2 cm above the objects, guiding the robot on where to navigate next.
+is trained on the same dataset of 1,500 environmental camera images as ModeNet.
+The ground truth for NavNet involves waypoints defined as positions 2 cm above the objects, guiding the robot on where to move next.
 Like ModeNet, data augmentation techniques such as rotation and contrast adjustments were used to enhance training robustness.
 NavNet employs Mean Squared Error (MSE) loss and is optimized with the Adam optimizer.
-It achieved an impressive accuracy score of 0.93, indicating high effectiveness in predicting accurate navigation waypoints based on visual inputs.
+It achieved an impressive accuracy score of 0.93, indicating high effectiveness in predicting accurate motion-planning waypoints based on visual inputs.
 Fig.
-7 shows basic demo of NAVINACT in navigation mode.
-4.1.2 ModeNet and NavNet for realworld experiments
-For real-world experiments, ModeNet and NavNet were trained using a more extensive dataset, incorporating both wrist and environment camera images totaling 6,200.
-This approach enhanced the diversity and representativeness of the training data, ensuring robustness and accuracy in real-world applications.
+7 shows basic demo of PLANRL in motion-planning mode.
+4.1.2 ModeNet and NavNet for real-world experiments
+For real-world experiments, ModeNet and NavNet were trained using a more extensive dataset of 6200 images, incorporating both wrist and environment camera images.
+The dataset captures a wider range of object perspectives, lighting conditions, and spatial contexts.
+This expansion improves the robustness and accuracy in real-world applications.
 4.2 Simulation Experiments
 Our evaluation suite includes three tasks from MetaWorld [37], each using sparse 0/1 task completion rewards at the end of each episode.
-These tasks are a subset of those evaluated in IBRL [28] and cover medium and hard difficulty levels as categorized in [35].
+These tasks cover medium and hard difficulty levels as categorized in [35].
 The specific environments used are assembly-v2, box-close-v2, and coffee-push-v2.
@@
 Although these tasks are relatively simple, they are sufficient to differentiate between stronger methods, especially in terms of sample efficiency and generalizability.
-Implementation of NAVINACT and Baselines:
-NAVINACT employs Behavior Cloning (BC) for imitation learning (IL) and Twin Delayed Deep Deterministic Policy Gradient (TD3) for reinforcement learning (RL).
+Implementation of PLANRL and Baselines:
+PLANRL employs Behavior Cloning (BC) for imitation learning (IL) and Twin Delayed Deep Deterministic Policy Gradient (TD3) for reinforcement learning (RL).
 A ResNet-18 vision encoder is used for BC, and random-shift image augmentation is applied to enhance RL performance.
-NAVINACT is benchmarked against three baselines: IBRL [28], RL-MN (RL with ModeNet and NavNet), and standard RL.
+PLANRL is benchmarked against three baselines: IBRL [28], RL-MN (RL with ModeNet and NavNet), and standard RL.
 IBRL uses a pre-trained IL policy, $\mu_{\psi}$, to aid exploration during RL and to assist in target value estimation during Temporal Difference (TD) learning.
-Likewise, NAVINACT also uses IL policy within the InteractNet, that speeds up RL during object interaction.
+Likewise, PLANRL also uses IL policy within the InteractNet, that speeds up RL during object interaction.
 RL with ModeNet and NavNet, which we refer as RL-MN, incorporates ModeNet and NavNet to accelerate exploration by predicting waypoints, reducing the number of exploration steps needed in RL.
@@
 4.3 Results
-Following the training of NAVINACT and the baselines, we evaluate its performance on unseen environments and compare it against baseline methods.
+Following the training of PLANRL and the baselines, we evaluate its performance on unseen configuration of object positions in the environments and compare it against baseline methods.
 Performance is assessed using training and evaluation scores, which indicate sample efficiency and generalizability.
 Fig.
-8 illustrates the comparative results for the all the metaworld environment, including training and evaluation scores for various initial positions.
+8 illustrates the comparative results for all the metaworld environments, including training and evaluation scores for various initial positions.
 (a) Assembly Env
@@
 (c) CoffeePush Env
-(d) Training with random inital Objects position.
-(e) Training with random inital Objects position.
-(f) Training with random inital Objects position.
+(d) Training with random initial Objects position.
+(e) Training with random initial Objects position.
+(f) Training with random initial Objects position.
 (g) Evaluation with random initial Objects position.
@@
 (l) Evaluation with random initial Objects and gripper position.
-Figure 8: Performance Evaluation of NAVINACT Across MetaWorld Tasks.
-This figure presents the training and evaluation results for three challenging MetaWorld environments: Assembly, BoxClose, and CoffeePush.
-The top row depicts the environments in their initial configurations.
-The middle row shows the training performance of NAVINACT under conditions where object positions are randomized.
-The bottom two rows illustrate the evaluation performance under different randomized scenarios: varying initial object positions and combined randomization of both object and gripper positions.
-The results demonstrate NAVINACT’s robustness and adaptability to varying task complexities, maintaining high performance across different training and evaluation conditions.
-Both NAVINACT and IBRL utilize a trained BC policy, with NAVINACT achieving a higher evaluation score of 0.38 in the assembly environment, while RL-MN and RL do not incorporate this policy.
-Our evaluation is focused on 3 key metrics — training score for the simulation tasks with randomization with respect to objects position, evaluation score with randomization with respect to objects position and evaluation score with randomization with respect to random initial gripper (as shown in Fig.
-6 as well as object positions.
-The performance of NAVINACT, IBRL, RL-MN, and standard RL is compared across the Meta-World environments over 60k timesteps and 7 seeds.
-NAVINACT outperforms baselines across all tasks in terms of both sample efficiency and final performance, solving all tasks within 30K samples.
-4.3.1 Does NAVINACT Lead to Better Sample Efficiency?
-To assess sample efficiency, we compare the training curves of NAVINACT with those of IBRL, RL-MN, and RL across various MetaWorld tasks.
+Figure 8: Performance Evaluation of PLANRL Across MetaWorld Tasks.
+This figure presents the training and evaluation results for three challenging MetaWorld environments: (a) Assembly, (b) BoxClose, and (c) CoffeePush.
+The top row depicts these environments in their initial configurations.
+The middle row (d), (e) and (f) shows the training performance of PLANRL under conditions where object positions are randomized within environments (a), (b), and (c).
+The bottom two rows illustrate the evaluation performance under different randomized scenarios: varying initial object positions and combined randomization of both object and gripper positions for each environment.
+The results demonstrate PLANRL’s robustness and adaptability to varying task complexities, maintaining high performance across different training and evaluation conditions.
+Both PLANRL and IBRL utilize a trained BC policy, with PLANRL achieving a higher evaluation score of 0.38 in the assembly environment, while RL-MN and RL do not incorporate this policy.
+Our evaluation is focused on 3 key metrics — training score for the simulation tasks with randomization with respect to object position, evaluation score with randomization with respect to object position, and evaluation score with randomization with respect to random initial gripper (as shown in Fig.
+6) as well as object positions.
+The performance of PLANRL, IBRL, RL-MN, and standard RL is compared across the Meta-World environments over 60k timesteps and 7 seeds.
+PLANRL outperforms baselines across all tasks in terms of both sample efficiency and final performance, solving all tasks within 30K samples.
+4.3.1 Does PLANRL Lead to Better Sample Efficiency?
+To assess sample efficiency, we compare the training curves of PLANRL with those of IBRL, RL-MN, and RL across various MetaWorld tasks.
 These tasks represent different levels of complexity, categorized into medium, hard, and very hard tiers.
-NAVINACT demonstrates better sample efficiency, as evidenced by faster convergence and higher success rates with fewer interaction steps as shown in Fig.
+PLANRL demonstrates better sample efficiency, as evidenced by faster convergence and higher success rates with fewer interaction steps as shown in Fig.
 8(d),8(e) and 8(f).
-This improvement is attributed to NAVINACT’s mode prediction, which effectively reduces task complexity by narrowing down the action space at each step.
-NAVINACT consistently achieves an impressive success rate of 85±5% during training and learns quickly within 30k sample steps.
+This improvement is attributed to PLANRL’s mode prediction, which effectively reduces task complexity by narrowing down the action space at each step.
+PLANRL consistently achieves an impressive success rate of 85±5% during training and learns quickly within 30k sample steps.
 Figure 9: Experimental Setup
-4.3.2 Does NAVINACT Generalize to Unseen Conditions?
-To evaluate generalization, we test NAVINACT in conditions not included in the training data, such as variations in object positions, and initial gripper position.
-NAVINACT shows strong generalization capabilities, significantly outperforming IBRL and pure RL methods in unseen scenarios as indicated in Fig.
+4.3.2 Does PLANRL Generalize to Unseen Conditions?
+To evaluate generalization, we test PLANRL in conditions not included in the training data, such as variations in object positions, and initial gripper position.
+PLANRL shows strong generalization capabilities, significantly outperforming IBRL and pure RL methods in unseen scenarios as indicated in Fig.
 8(g),8(h),8(i),8(j),8(k) and 8(l).
 The mode prediction network plays a crucial role in adapting to new conditions by dynamically adjusting the operational mode, which is not achievable by methods solely relying on IL.
-Our algorithm maintains a nearly equivalent rate of 80±5% in evaluation phases across unseen environments.
-NAVINACT consistently outperforms other baselines in MetaWorld tasks, demonstrating its robustness to variability and changes in the environment.
-We see that, NAVINACT matches or surpasses other baselines in Meta-World.
-The results indicate that, NAVINACT outperforms the other three methods in both training and evaluation.
-Notably, as randomization increases, the performance of IBRL, RL-MN, and standard RL decreases, while NAVINACT maintains consistent performance and demonstrates the robustness to variability and changes in environment.
+Our algorithm maintains a nearly equivalent rate of 80±5% in evaluation phases across unseen configurations of object position in the environments.
+We see that, PLANRL matches or surpasses other baselines in Meta-World.
+Notably, as randomization increases, the performance of IBRL, RL-MN, and standard RL decreases, while PLANRL maintains consistent performance and demonstrates robustness to variability and changes in the environment.
 4.4 Real-world Experimental Setup
@@
 We detect whether the gripper is holding the block by checking if the gripper width is static and the desired gripper width is smaller than the actual gripper width.
-The success detector returns 1 if the end effector has move upward by at least 2cm while holding the block.
+The success detector returns 1 if the end effector has moved upward by at least 2cm while holding the block.
 The maximum episode length is 120.
@@
 The initial location of the soft toy and the block is in a fixed position.
-The soft toy is initialized so that its visible from the wrsit camera view, see Fig.10(c).
+The soft toy is initialized so that it is visible from the wrist camera view, see Fig.10(c).
 We use 20 prior demonstrations.
-This task uses environment camera images as observations as well because the wrist-camera loses sight of the block after picking up the soft-toy.
-We detect whether the gripper is holding the oft-toy by checking if the gripper width is static and the desired gripper width is smaller than the actual gripper width.
-The success returns 1 when the gripper width is static, the end-effector position are near the threshold of desired block position.
-The maximum episode legnth is 300.
+This task uses environment camera images as observations as well because the wrist-camera loses sight of the block after picking up the soft toy.
+We detect whether the gripper is holding the soft toy by checking if the gripper width is static and the desired gripper width is smaller than the actual gripper width.
+The success returns 1 when the gripper width is static, the end-effector position is near the threshold of desired block position.
+The maximum episode length is 300.
 (a) Lift
@@
 Bottom: Performance evaluation on the lift task under both fixed and randomized conditions.
-The training curves depict the success rate over time, demonstrating NAVINACT’s effective learning and adaptability in varying scenarios.
-As the primary goal of our real-world evaluations is to compare sample-efficiency and performance of various algorithms, we design rule-based success detectors and perform manual reset between episodes to ensure accurate reward and initial conditions.
-Note that sparse 0/1 reward from the success detector is the only source of reward.
+The training curves depict the success rate over time, demonstrating PLANRL’s effective learning and adaptability in varying scenarios.
+As the primary goal of our real-world evaluations is to compare sample efficiency and performance of various algorithms, we design rule-based success detectors and perform manual reset between episodes to ensure accurate reward and initial conditions.
+Note that the sparse 0/1 reward from the success detector is the only source of reward.
 4.5 Real-world Experiment Results
-Does NAVINACT Work Effectively in Real-World Scenarios?
+Does PLANRL Work Effectively in Real-World Scenarios?
 Fig.
-10 presents the training curves of NAVINACT compared to IBRL and across all different tasks.
-Each task is allowed a different interaction budget based on its difficulty, and the success rate is measured over each 1000-step interval during training.
-NAVINACT consistently learns faster than IBRL, and in the most challenging task, Pick-and-Place, it is the only method that surpasses behavior cloning (BC) under a 5K interaction budget.
+10 presents the training curves of PLANRL compared to IBRL and across all different tasks.
+Each task is allowed different interactions based on its difficulty, and the success rate is measured over each 1000-step interval during training.
+PLANRL consistently learns faster than IBRL, and in the challenging task, Pick-and-Place, it is the only method that surpasses behavior cloning (BC) under 5K interactions.
 In real-world evaluations, we assessed the final performance of each method using identical initial conditions.
-In the Lift task, NAVINACT achieves the highest success rate in both regular and “Hard Eval” settings, where the block’s initial position is more challenging.
-Even when the task difficulty increases, NAVINACT maintains a high success of 90% rate, demonstrating its robustness to distributional shifts and unseen scenarios, unlike BC, which fails in harder settings due to its reliance on the static training dataset as shown in Fig.
+In the Lift task, PLANRL achieves the highest success rate in both regular and “Hard Eval” settings, where the block’s initial position is more challenging.
+Even when the task difficulty increases, PLANRL maintains a high success of 90% rate, demonstrating its robustness to distributional shifts and unseen scenarios, unlike BC, which fails in harder settings due to its reliance on the static training dataset as shown in Fig.
 10.
-How does NAVINACT perform for long-horizon tasks?
-The Pick-and-Place task, the most difficult due to two stage task manipulation, further highlights NAVINACT’s advantages as displayed in Fig.
+How does PLANRL perform for long-horizon tasks?
+The Pick-and-Place task, the most difficult due to two-stage task manipulation, further highlights PLANRL’s advantages as displayed in Fig.
 10.
-While BC also doesn’t performs well due to suboptimal demostration, other method IBRL fail to match its success rate.
-While BC achieves a limited success rate of 20%, NAVINACT surpasses this by reaching a 75% success rate, even in complex scenarios where small errors can lead to significant failures.
+While BC also doesn’t perform well due to suboptimal demonstration, other methods IBRL fail to match its success rate.
+While BC achieves a limited success rate of 20%, PLANRL surpasses this by reaching a 75% success rate, even in complex scenarios where small errors can lead to significant failures.
 5 Discussion
 5.1 Conclusion
-In this paper, we presented NAVINACT, a hierarchical policy framework that combines NavNet for waypoint planning, ModeNet for dynamic mode switching, and InteractNet for fine-grained manipulation.
-By integrating imitation learning (IL) with reinforcement learning (RL), NAVINACT effectively addresses challenges of sample efficiency and distributional shift, outperforming existing methods.
-Through extensive simulations and real-world experiments, we demonstrated NAVINACT’s superior performance, generalizability, and adaptability across various tasks.
-Our approach learns faster and requires fewer interaction steps due to the innovative integration of mode and waypoint prediction, making it a scalable and robust solution for complex robotic manipulation in real-world scenarios.
+In this paper, we presented PLANRL, a hierarchical policy framework that combines NavNet for waypoint motion-planning, ModeNet for dynamic mode switching, and InteractNet for fine-grained manipulation.
+By integrating imitation learning (IL) with reinforcement learning (RL), PLANRL effectively addresses challenges of sample efficiency and distributional shift, outperforming existing methods.
+Through extensive simulations and real-world experiments, we demonstrated PLANRL’s superior performance, generalizability, and adaptability across various tasks.
+Our approach learns faster and requires fewer interaction steps due to the integration of mode and waypoint prediction, making it a scalable and robust solution for complex robotic manipulation in real-world scenarios.
 5.2 Limitation and Future Work
-While NAVINACT demonstrates significant advancements in sample efficiency and generalizability, one limitation of our approach is the reliance on collecting high-quality demonstration data, which can be time-consuming and expensive.
-This dependence on expert-generated data constrains the scalability and applicability of our framework across a wider range of tasks and environments.
-Future work could focus on developing methods that allow non-experts to collect usable data quickly and efficiently, potentially through resources such as videos in the wild.
+While PLANRL demonstrates significant advancements in sample efficiency and generalizability, one limitation of our approach is the reliance on collecting high-quality demonstration data, which can be time-consuming and expensive.
+Future work could focus on developing methods that allow non-experts to collect usable data quickly and efficiently, potentially through resources such as videos in the wild or collecting demonstrations using sketches.